OpenStudy (anonymous):

How do you factor 8z^3-75z=0?

OpenStudy (anonymous):

pull out a z and you'll get a squared function

OpenStudy (anonymous):

then use quadratic formula if you'd lik

OpenStudy (anonymous):

Factorising is when you take a common term out of the equation and use it to make the equation simpler. In your case, the common term we can take out is z, but also whatever is a factor of both 8 and 75.

OpenStudy (anonymous):

\[8z ^{3}-75z=0\] \[z(8z ^{2}-75)=0\] \[z(2z \sqrt{2}+5\sqrt{3})(2z \sqrt{2}-5\sqrt{3})=0\]